OpenStudy (anonymous):

How do I turn the equation y=2x + 9 into standard form

OpenStudy (anonymous):

ello :D

OpenStudy (anonymous):

do you know the standard equation?

OpenStudy (anonymous):

i CAN HELP YOU @mammania

OpenStudy (anonymous):

can u help me 11. (03.06 LC) Choose the equation below that represents the line that passes through the point (-7, 2) and has a slope of −3. (5 points) y − 2 = −3(x + 7) y + 7 = −3(x − 2) y + 2 = −3(x − 7) y − 7 = −3(x + 2) Score: 0 of 5

OpenStudy (anonymous):

could u help me @atreyu6s3x6

OpenStudy (anonymous):

okay so do you know the point slope equation? @Jesuslover14

OpenStudy (anonymous):

sorry if i reply slow im helping a buddy with some work also :) @Jesuslover14

OpenStudy (anonymous):

yes is it y=mx+b

OpenStudy (anonymous):

that is slope intercept, point slope is y-y1=m(x-x1)

OpenStudy (anonymous):

so we know that: m=-3 and it passes through point: (-7,2) so simply plug in the numbers @Jesuslover14

OpenStudy (anonymous):

SORRY ALREADY HAVE IT

Directrix (directrix):

@Jesuslover14 Create your own threads with your questions. Do not be a "thread-buster." @atreyu6s3x6 Please answer OP @mammamia_1257 's question and avoid the distraction of the thread buster. Thanks.

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

dude shes not here i was waiting for her to answer me so i could walk her through it then he asked for help. @Directrix

OpenStudy (anonymous):

so chill @Directrix

OpenStudy (anonymous):

i know that you can turn the equation into standard form from just this. You don't need Point Slope form to change it. @atreyu6s3x6 and Sorry for not responding I was trying to finish the rest of my geometry homework

OpenStudy (anonymous):

yea everything after jesuslover is not directed at you and its quite alright :) @mammamia_1257

Directrix (directrix):

Standard Form: the standard form of a line is in the form Ax + By = C where A is a positive integer, and B, and C are integers. @mammamia_1257

Directrix (directrix):

@mammamia_1257 How do I turn the equation into standard form y=2x + 9 Subtract 2x from both side of the equation y - 2x = 2x - 2x + 9 y - 2x = 9 Put the x term first on the left side -2x + y = 9 Multiply each term by -1 to complete standard form. 2x - y = - 9 Standard Form

OpenStudy (anonymous):

Thank you @Directrix I ended up figuring it out, I've been having mind blanks lately. Thank you.

Directrix (directrix):

You are welcome.